BCAL Reconstruction Meeting 2015-01-08

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m. EDT

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204


  1. BCAL Reconstruction Issues
  2. BCAL Reconstruction Algorithms
  3. Will's reconstructed energy info

Action Items

  1. Ask Richard Jones about Truth info from BCAL hadronic showers (Zisis, Will L., Justin)
  2. Shower curvature tables: compare KLOE to IU codes: check gap in cluster handling between codes (Tegan, Andrei)
  3. z-coordinate determination from up/down amplitude ratio (Tegan, Andrei)
  4. Cleanup: (Tegan, Will L.)
    1. Set variable “failures” to a number other than zero (e.g. -9999 or similar)
    2. Comb code for constants to be pulled out into a database
  5. Measure data pedestals using LV=0 setting (Mark)
  6. Sampling fraction tables (Tegan)


  1. Announcements
  2. Action Items
  3. Review and pruning BCAL Reconstruction Issues
  4. Commissioning
  5. Calibration
    1. Calibration Constants (Elton)
    2. Cosmics data analysis update (Tegan)
  6. Simulations
    1. BDT effort status quo & plans - (Zisis, Tegan)
  7. Any other business


Attendees: Will L., Will M. (CMU), Elton, Mark D., David (JLab), Christina, George (Athens), Tegan, Ahmed, Zisis (Regina)

  1. Announcements:
    1. Zisis reported that Justin Stevens will visit Regina next week to work on BDT.
    2. Elton reported on the accelerator schedule. The accelerator will start Feb 27 and will have priority in March, with a goal to deliver to all end stations. We do not know if Hall D will be locked up for all of March. Beam will be delivered to us on Apr 2 with priority to physics (us). It is not in our best interest to try to accept beam in March as there is much work to do on hardware and software, in order to be in an optimal position for the April running. As a result, we decided to take BCAL cosmics in February and Noemi will go to JLab for this.
  2. Action Items: no updates.
  3. Review and pruning BCAL Reconstruction Issues:
    1. General comments: We went through the list. A few items have been done. A few have been forgotten and Zisis will try to figure out from past meetings what they were. Mark and Tegan are working on a few. New simulations will be examined to see which of these issues persist and also organize the tasks based on priority.
    2. Specific comments:
      1. Sims/FPGA: done; poor resolution: Tegan will examine.
      2. Dark hits: David sent an email during the meeting, reminding us of the status quo. Briefly, the simulation overestimates their amplitude compared to real data. It turns out David did not check in any change regarding this, but provided the e-mail showing what was done in October.
      3. Max ADC 12 bits: this was discussed because NorthWestern wanted to look at J/psi->e+e- and wanted to check whether ADC would saturate. Nth was done. Zisis will contact Shaun about this. Elton suggested that we handle pulses from the data (using appropriate conversion factors, integral/peak, flash dynamic range and nominal coefficients in ccdb). The pulse handling can be done on the fly in MCSmear.
      4. Sampling fluctuations means Andrei's sampling fraction tables, listed in our Action Items.
      5. Attenuation length. We want to pull out constants such as this one from the code and get them from ccdb (see Action Item). We need to arrive at using the same calibration constants in data as well as MC. Use Elton's Low-Level Coefficients table.
      6. Time-walk: fADC perhaps requires small t-w correction. The comment here is meant for TDCs and we are a long ways from using these. Can we use amplitude ratio for z-position? Andrei can comment.
      7. DBCALPoint: Mark is working on aspects of this. We have attenuation length values, and soon good values for the gain.
      8. Clustering: we only use ADCs in the real data until now. Splitting clusters based on time has been implemented but we need to check that it works. The issue at present is mainly that not enough energy is included into the cluster rather than too much. The split could become important at high rates.
      9. Match with tracks: Paul and Simon have worked on this, for FCAL triggers (BCAL and FCAL-BCAL triggers will follow). Will M. used his own routines to match tracks as the BCAL hits come before the Drict Chambers and so the default routines cannot be used to match tracks. The correct timing offsets for BCAL need to go into ccdb. Mark is working on this for pi0s, propagating them back to the target in a straight line using the speed of light. Longer term: compare method to one using charged particles.
  4. Commissioning: Tegan reported that the pedestal-check script is working, and it produces PDF plots versus run number and the sigma from the fit. He is now working on a skim script/plugin to put pedestals in a file. Elton commented that we need to track the pedestal stability over time, look for periodicity, how much it varies, outliers, etc. Things look basically good, but we should investigate these things.
  5. Calibration
    1. Elton reported on the updates and corrections to the Low-Level Calibrations document. The idea is to have these implemented in the code but taken from the ccdb. Zisis will contact Shaun to see if he can implement Table 1 (page 6) into the ccdb now. Please read the document and give feedback.
    2. Cosmics data analysis update (Tegan).
    3. Tegan asked for the Trigger Paddle location. Marj=k said thet are in Roc 37, Slot 19. First 4 channels are the bottom trigger counters, next two are the top ones and Will M. confirmed.
  6. Simulations
    1. BDT effort status quo & plans: Zisis reported that during Justins's visit we will look at improving the BDT to reject bad clusters, investigate new variables for the BDT, look at a regression BDT to improve the energy resolution. In preparation we have pin-pointed the appropriate revision for this work (rev. 16089) and Tegan is running single photon, single pi+/-, single pi0, bggen MC data sets both with and without dark hits. Longer term: we will discuss how the methods we're developing can be translated to real data once we reach that point.
    2. Grid efforts in Regina (Canadian WestGrid Cluster). Mark Ito is planning to run DC3 at JLab only and investigate getting date off tape (logistics, rates, etc). Zisis will work with Richard Jones to tie WestGrid into our next big Data Challenge.
  7. Any other business:
    1. Elton reported on his investigation on light reflection in the BCAL. His model shows that the ratio of integral to peak is not constant as a function of position, with a minimum at 1/4 along the length. Main features of the data are reproduced by the model, namely larger attenuation length than expected from Regina test-best measurements with cosmics. Christina asked about the LED reflections and Elton answered that those are not applicable to the simulation. The model results in a 10% reflection coefficients. Elton asked for folks to read this document and send feedback.